One positive number is one-fifth of another number. The difference between the two numbers is 244, find the numbers.
Answer by edjones(8007)

x=y/5
y-x=244
.
y - y/5 = 244 Substitute y/5 for x.
5y-y=1220 multiply each side by 5.
4y=1220
y=305
x=61
.
Ed